OpenStudy (anonymous):

The plates of a parallel plate capacitor are quadratic with side length l. When the plates are separated by a 1.15 cm thick sheet of pyrex glass, a capacitance of C = 17.5 pF is measured. What is the length l ?

OpenStudy (anonymous):

c=(k(epslon not)*A)/d here k=dielectric constant of pyrex glass=k=4.6 A=l*l,d=1.15 cm,c=17.5pF.put these value in above formula and find the value of l....